# NutriGen: Personalized Meal Plan Generator Leveraging Large Language Models to Enhance Dietary and Nutritional Adherence

### v1

@@ I INTRODUCTION
The rise of machine learning (ML) and data analysis techniques has further enhanced the effectiveness of personalized dietary recommendations [7]. By leveraging user data, these advanced systems can generate insights into optimal eating patterns while considering individual preferences. For example, DietOS employs optimization algorithms to develop customized diet plans tailored to the management of chronic diseases [8]. Similarly, IoMT-assisted systems, such as the approach by Iwendi et al., utilize ML to create precise, patient-specific dietary recommendations, excelling in clinical environments [9]. ChatDiet integrates large language models (LLM) to deliver interactive and engaging dietary advice via chatbot interfaces, significantly enhancing user interaction [6]. Additionally, optimization techniques like Particle Swarm Optimization (PSO) have been applied to generate meal plans that efficiently balance nutritional goals [10]. These innovative methodologies illustrate the potential of computational tools in addressing dietary challenges.
@@
However, despite their potential, existing dietary recommendation systems face notable limitations that restrict their practical application. Methods such as PSO, while efficient at balancing nutritional objectives, often fail to account for user-specific preferences like cultural restrictions or diverse dietary needs, which reduces their real-world usability [10]. Similarly, IoMT-assisted systems, though highly precise and robust in clinical settings, are not well-suited for the general population seeking practical, everyday dietary guidance. Their reliance on continuous medical monitoring, clinical validation, and specialized sensors limits their usability outside structured healthcare environments, reducing their practicality for everyday meal planning [9]. ChatDiet excels in user engagement but struggles with scalability due to high computational costs, platform dependencies, and reliance on continuous user input [6]. Moreover, many existing systems focus solely on recommending individual food items rather than offering comprehensive and cohesive diet menus. This leaves users with the burden of assembling balanced meal plans on their own. Additionally, these systems frequently lack features such as alternative food suggestions, consideration of ingredient availability, or integrated recipes and cooking instructions. As a result, users may find it challenging to implement the recommendations in their daily lives. These challenges often lead to steady declines in engagement over time, with many participants discontinuing their use of the tools altogether. Addressing these gaps requires the development of adaptable, user-centered frameworks that provide structured, actionable, and personalized dietary guidance [11, 12, 13].
@@ II-C Prompt Engineering
To further enhance reliability, we incorporated few-shot prompting by providing structured examples within the prompt. These examples serve as reference points, guiding the LLM to follow a consistent format while allowing diversity in meal recommendations. Additionally, our prompt design supports flexible constraints, enabling users to prioritize specific aspects such as high-protein meals or low-sugar alternatives. Below is the core section of the structured template prompt used in all our tests. The complete prompt and implementation code will be available for reference.
@@ III-B Selected Models
  • Llama 3.1-70B: Developed by Meta, this model comprises 70 billion parameters and has been trained on approximately 15 trillion tokens, demonstrating enhanced performance in language understanding and generation tasks.
  • Llama 3.1-8B: A smaller variant from Meta, this 8 billion parameter model is designed for efficiency while maintaining robust language processing abilities.
@@
  • DeepSeek-V3: Released by DeepSeek in December 2024, this model features 671 billion active parameters and a context length of 128K tokens, excelling in multilingual tasks, particularly in English and Chinese.
  • Gemini 2.0 Flash Exp: Developed by Google DeepMind, this model integrates advanced language understanding with reinforcement learning techniques, aiming for improved reasoning and problem-solving abilities.
  • Gemini 1.5 Pro: Offers advanced language understanding and reasoning capabilities, optimized for professional application.
  • GPT-4o: OpenAI’s latest model supports both text and image inputs, maintaining high intelligence levels across various tasks.
  • GPT-4o Mini: A compact version engineered to deliver high performance in a more resource-efficient package.
  • GPT-3.5 Turbo: An optimized version of GPT-3, offering faster response times and improved efficiency in language tasks.
@@ IV-A Processing Time Comparison
DeepSeek V3 138.09 96.6 88 - - - - 100 82.84 113
Gemini 1.5 Pro 588 173.59 87.75 153.5 445.5 75 76.75 96 83 52.59
Gemini 2 Flash 589.33 173.59 87.75 153.5 445.5 75 103.41 96 83 48
GPT-3.5 Turbo 138.09 73.59 37.75 53.5 45.5 25 26.75 46 32.84 62.59
GPT-4o 50 24 38 45 563 41.6 50 36 74.6 52.59
@@
Llama 3.1 8B 88.09 58.33 50 60 46 58.33 26.75 - 50 43.33
Llama 3.1 70B 45 60 50 50 46.66 40 48.33 46 50 63
To evaluate the computational efficiency of each model, we measured the total processing time required to generate 10 outputs. The results, presented in Fig.  2, show significant variations in execution time across different models. As expected, smaller and optimized models such as GPT-3.5 Turbo, Gemini 2.0 Flash, and Claude 3.5 Haiku demonstrated the fastest processing times, completing the task in significantly less time compared to larger models.
@@ IV-C Adherence of Meal Plans to User-Specified Targets
DeepSeek V3 190.61 4.85
Gemini 1.5 Pro 182.16 10.44
Gemini 2 Flash 179.16 9.74
GPT-3.5 Turbo 54.16 3.68
GPT-4o 189.76 13.47
@@ IV-D Evaluation of Meal Plan Quality and Completeness
V CONCLUSIONS
@@
Despite its strengths, NutriGen has certain limitations that warrant further improvement. Some language models struggled with output token constraints, leading to incomplete meal plans that omitted key nutritional details. Additionally, calorie estimation discrepancies across models highlight the need for a more robust retrieval mechanism to ensure greater consistency and accuracy in nutritional data generation.
To address these challenges, future work will explore RAG to dynamically fetch real-time USDA data, minimizing retrieval errors and refining nutrient calculations. Moreover, NutriGen currently lacks interactivity, limiting user control over meal plan adjustments. Integrating a ChatBot interface would allow users to modify meal plans, adjust dietary restrictions, and update their nutrition profiles in real time, enhancing personalization and user engagement. Expanding NutriGen to leverage multimodal LLMs with food image recognition could further improve meal recommendations by incorporating visual dietary analysis. This capability would enable the system to assess ingredient availability, suggest alternative meal options, and enhance dietary adherence through a more comprehensive and interactive approach to personalized nutrition planning.

### v2

@@ I INTRODUCTION
The rise of machine learning (ML) and data analysis techniques has further enhanced the effectiveness of personalized dietary recommendations [7, 8, 9]. By leveraging user data, these advanced systems can generate insights into optimal eating patterns while considering individual preferences. For example, DietOS employs optimization algorithms to develop customized diet plans tailored to the management of chronic diseases [10]. Similarly, IoMT-assisted systems, such as the approach by Iwendi et al., utilize ML to create precise, patient-specific dietary recommendations, excelling in clinical environments [11]. ChatDiet integrates large language models (LLM) to deliver interactive and engaging dietary advice via chatbot interfaces, significantly enhancing user interaction [6]. Additionally, optimization techniques like Particle Swarm Optimization (PSO) have been applied to generate meal plans that efficiently balance nutritional goals [12]. These innovative methodologies illustrate the potential of computational tools in addressing dietary challenges.
@@
However, despite their potential, existing dietary recommendation systems face notable limitations that restrict their practical application. Methods such as PSO, while efficient at balancing nutritional objectives, often fail to account for user-specific preferences like cultural restrictions or diverse dietary needs, which reduces their real-world usability [12]. Similarly, IoMT-assisted systems, though highly precise and robust in clinical settings, are not well-suited for the general population seeking practical, everyday dietary guidance. Their reliance on continuous medical monitoring, clinical validation, and specialized sensors limits their usability outside structured healthcare environments, reducing their practicality for everyday meal planning [11]. ChatDiet excels in user engagement but struggles with scalability due to high computational costs, platform dependencies, and reliance on continuous user input [6]. Moreover, many existing systems focus solely on recommending individual food items rather than offering comprehensive and cohesive diet menus. This leaves users with the burden of assembling balanced meal plans on their own. Additionally, these systems frequently lack features such as alternative food suggestions, consideration of ingredient availability, or integrated recipes and cooking instructions. As a result, users may find it challenging to implement the recommendations in their daily lives. These challenges often lead to steady declines in engagement over time, with many participants discontinuing their use of the tools altogether. Addressing these gaps requires the development of adaptable, user-centered frameworks that provide structured, actionable, and personalized dietary guidance [13, 14, 15, 16].
@@ II-C Prompt Engineering
To further enhance reliability, we incorporated few-shot prompting by providing structured examples within the prompt. These examples serve as reference points, guiding the LLM to follow a consistent format while allowing diversity in meal recommendations. Additionally, our prompt design supports flexible constraints, enabling users to prioritize specific aspects such as high-protein meals or low-sugar alternatives. Below is the core section of the structured template prompt used in all our tests. The complete prompt and implementation code  [^2] https://github.com/SamanKhamesian/NutriGen is available for reference.
@@ III-B Selected Models
  • Llama 3.1-70B [21]: Developed by Meta, this model comprises 70 billion parameters and has been trained on approximately 15 trillion tokens, demonstrating enhanced performance in language understanding and generation tasks.
  • Llama 3.1-8B [21]: A smaller variant from Meta, this 8 billion parameter model is designed for efficiency while maintaining robust language processing abilities.
@@
  • DeepSeek-V3 [22]: Released by DeepSeek in December 2024, this model features 671 billion active parameters and a context length of 128K tokens, excelling in multilingual tasks, particularly in English and Chinese.
  • Gemini 2.0 Flash Exp [23]: Developed by Google DeepMind, this model combines language understanding with reinforcement learning to enhance reasoning and problem-solving.
  • Gemini 1.5 Pro [23]: Offers advanced language understanding and reasoning capabilities, optimized for professional application.
  • GPT-4o [24]: OpenAI’s latest model supports both text and image inputs, maintaining high intelligence levels across various tasks.
  • GPT-4o Mini: A smaller model designed for high performance with lower resource usage.
  • GPT-3.5 Turbo [25]: An optimized version of GPT-3, offering faster response times and improved efficiency in language tasks.
@@ IV-A Processing Time Comparison
DeepSeek V3 [22] 138.09 96.6 88 - - - - 100 82.84 113
Gemini 1.5 Pro [23] 588 173.59 87.75 153.5 445.5 75 76.75 96 83 52.59
Gemini 2 Flash [23] 589.33 173.59 87.75 153.5 445.5 75 103.41 96 83 48
GPT-3.5 Turbo [25] 138.09 73.59 37.75 53.5 45.5 25 26.75 46 32.84 62.59
GPT-4o [24] 50 24 38 45 563 41.6 50 36 74.6 52.59
@@
Llama 3.1 8B [21] 88.09 58.33 50 60 46 58.33 26.75 - 50 43.33
Llama 3.1 70B [21] 45 60 50 50 46.66 40 48.33 46 50 63
To evaluate the computational efficiency of each model, we measured the total processing time required to generate 10 outputs. The results, presented in Fig. 2, show significant variations in execution time across different models. As expected, smaller and optimized models such as GPT-3.5 Turbo, Gemini 2.0 Flash, and Claude 3.5 Haiku demonstrated the fastest processing times, completing the task in significantly less time compared to larger models.
@@ IV-C Adherence of Meal Plans to User-Specified Targets
DeepSeek V3 [22] 190.61 4.85
Gemini 1.5 Pro [23] 182.16 10.44
Gemini 2 Flash [23] 179.16 9.74
GPT-3.5 Turbo [25] 54.16 3.68
GPT-4o [24] 189.76 13.47
@@ IV-D Evaluation of Meal Plan Quality and Completeness
V FUTURE WORK
Despite its strengths, NutriGen has limitations that require further improvement. Some language models encountered output token constraints, leading to incomplete meal plans missing key nutritional details. Inconsistencies in calorie estimation also highlight the need for a more robust retrieval mechanism. Additionally, the use of simulated user data may limit the generalizability of the findings.
To address these issues, future work will explore RAG to dynamically fetch real-time USDA data, reducing retrieval errors and improving nutrient calculations. NutriGen also lacks interactivity, limiting user control over meal plan customization. Integrating a ChatBot interface would allow real-time updates to preferences, restrictions, and nutrition profiles, enhancing engagement and personalization. Further, incorporating multimodal LLMs with food image recognition could improve recommendations by analyzing visual dietary input, assessing ingredient availability, and suggesting alternatives. Although currently limited to English, NutriGen’s language-agnostic design allows adaptation to other languages through localized databases and multilingual LLMs like GPT-3.5 Turbo.
VI CONCLUSIONS
@@
VII ACKNOWLEDGMENT
This work was supported in part by the National Science Foundation under grant IIS-2402650. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding organizations.